Sa of 10 > Suppose a biologist studying the mechanical limitations of growth among different species of tulips monitors a nation preserve. He collects data on the heights of 10 different types of tulips in the reserve and rounds cach height to the nea centimeter. 25, 21,26, 24, 27,30, 29,25, 16,23 Compute the first quartile (Q1). the third quartile (Os), and the interquartile range (IQR) of the data set. 021 cm 03 28 cm IQR 25 cm
Answer #1

The quartiles are computed using the table with the data in ascending order For Q1 we have to compute the following position:

So, we have :

Q1 = 22.5 cm

327.5 cm

Now, we use the above values to calculate Inter-Quartile Range:

IQR = Q3-Q1 = 27.5-22.5 = 5 cm